OpenStudy (anonymous):

Solve using the principle of zero products (11a+7)(4a-12)=0

jimthompson5910 (jim_thompson5910):

(11a+7)(4a-12)=0 11a+7=0 or 4a-12=0 11a=-7 or 4a=12 a=-7/11 or a=12/4 a=-7/11 or a=3

OpenStudy (anonymous):

ok